OpenStudy (anonymous):

In an arthimetic sequence, the 8th term minus the 1st term equals 14. The sum of the 1st and 8th term is -2. Find the 4th term

jhonyy9 (jhonyy9):

1st term let a and the 8th term let h so than we know that h-a=14 h+a=-2 -------- 2h=12 h=6 6-a=14 6-14=a -8=a a=-8 h=6 the 4th term will be the 8th minus 1st so than -8+6=-2 so than the 4th term is -2

OpenStudy (anonymous):

Thank you so so much!

jhonyy9 (jhonyy9):

was my pleasure yw good luck bye